# Inverse Kinematics for Neuro-Robotic Grasping with Humanoid Embodied Agents

### v1

@@ Abstract
This paper introduces a novel zero-shot motion planning method that allows users to quickly design smooth robot motions in Cartesian space. A Bézier curve-based Cartesian plan is transformed into a joint space trajectory by our neuro-inspired inverse kinematics (IK) method CycleIK, for which we enable platform independence by scaling it to arbitrary robot designs. The motion planner is evaluated on the physical hardware of the two humanoid robots NICO and NICOL in a human-in-the-loop grasping scenario. Our method is deployed with an embodied agent that is a large language model (LLM) at its core. We generalize the embodied agent, that was introduced for NICOL, to also be embodied by NICO. The agent can execute a discrete set of physical actions and allows the user to verbally instruct various different robots. We contribute a grasping primitive to its action space that allows for precise manipulation of household objects. The new CycleIK [^2] Available on GitHub: https://github.com/jangerritha/cycleik method is compared to popular numerical IK solvers and state-of-the-art neural IK methods in simulation and is shown to be competitive with or outperform all evaluated methods when the algorithm runtime is very short. The grasping primitive is evaluated on both NICOL and NICO robots with a reported grasp success of 72% to 82% for each robot, respectively.
@@ I Introduction
Precise inverse kinematics is the key factor for successful grasping motions in many robotic manipulation applications. IK methods predict joint angles that align the robot end effector with a given position and orientation. They therefore collapse the action space of higher layered motion planners to the 6D pose of the end effector, while the alternative is to control the robot directly in its joint space which poorly scales for robots with varying degrees of freedom (DoF). The complexity of the IK task increases drastically for redundant kinematic chains with more than six DoF, since for non-redundant chains a specific pose maps exactly to one joint configuration, while for redundant manipulators infinite solutions generally exist, which makes the problem ambiguous. The number of neuro-inspired IK methods that are precise enough to be applied to physical hardware is increasing but there are only a few approaches that utilize them for neural motion planning[1, 2]. Many motion planners exist that use neural networks to sample the robot joint space and deploy classical planning algorithms like MPNet[3]. Thus, only the runtime of the sampling decreases, while the latency for the motion planning itself stays high. Other approaches plan directly in Cartesian space and transform the plan to the joint space via inverse kinematics such as CppFlow[1]. Neural networks that learn to solve the IK task, which allows for parallel batched queries, can dramatically accelerate Cartesian planning methods but decrease the precision. Numerical IK approaches have to solve each pose individually, resulting in a trade-off between precision and runtime, for which we propose one of the most runtime-efficient motion planning systems available.
@@
The neural inverse kinematics solver CycleIK[4] is the central component of our proposed motion planning approach. CycleIK is scaled to arbitrary robot designs by a replacement of the loss functions for the positional and rotational error with two new metrics that are derived from the Smooth L1 loss[5]. The low-level motion planner interpolates a Bézier curve between the current pose and a target pose and then converts the Cartesian plan into a joint space trajectory in a single step via CycleIK. Our method is deployed with the high-level embodied agent that is introduced to the NICOL[6] robot by Allgeuer et al.[7]. We generalize the embodied agent architecture over our humanoid robot platform to also be embodied the NICO[8] robot. The adult-sized NICOL robot is capable of manipulating realistic household objects, while the child-sized NICO can only manipulate toy-sized objects.
@@
The new CycleIK methodology is evaluated against two well-established IK methods available in Moveit[12], the genetic algorithm BioIK[13] and the Jacobian-based least squares optimizer Trac-IK[14]. The algorithms are compared on five different robot platforms: NICO, NICOL, Franka Emika Panda[15], NASA Valkyrie[16], and the Fetch[17] robot. In addition, the proposed IK method is compared against two state-of-the-art neuro-generative IK approaches by Limoyo et al.[18] and Ames et al.[19]. Finally, the grasp success of our method is evaluated on the NICO and the NICOL robots, for $100$ grasp attempts on each robot.
@@ III Method
The presented method fundamentally updates our neuro-inspired IK approach CycleIK[4] to enable platform independence for our method. Most importantly, the loss functions for the positional and rotational errors are exchanged by new Smooth L1-based metrics. The CycleIK multi-layer perceptron (MLP) is a single-solution IK solver that deterministically predicts a single vector of joint angles for a specific pose. Secondary constraints can be applied during the training to influence the kinematic behavior of the model. For non-redundant robotic chains with up to six degrees of freedom, only the MLP network is applicable, as no null space exists. CycleIK also provides a generative adversarial network that is capable of sampling the null space of a specific pose for redundant kinematic chains with higher DoF. However, as this study does not consider obstacle-avoiding motion planning, the MLP is sufficient since it is more applicable and will therefore be used to solve the inverse kinematics problem in our experiments.
@@
Figure 3: Overview of the CycleIK training algorithm. Kinematic constraints can either be set in the joint space or Cartesian space.
@@ III-B Training
We utilize the geometric forward kinematics solver PyTorch Kinematics[30] that allows for batched FK queries as it calculates the end effector pose for a particular joint state by matrice multiplications and allows for computational acceleration by utilizing the GPU via PyTorch. The operations used by PyTorch Kinematics are differentiable and can therefore be used during the network training without blocking the propagation of the gradient. Every training iteration processes a batch of poses with a minimum size of $100$, while the exact batch size is a hyperparameter that is optimized per robot. The learning rate is decreased to zero in linear steps after each training epoch.
@@ III-C Metrics
$l(a,b)=\begin{cases}\tfrac{1}{2}(a-b)^{2}/\beta,&\text{if $|a-b|<\beta$}\\ |a-b|-\tfrac{1}{2}*\beta,&\text{otherwise}\end{cases}$ (5)
@@
First, the element-wise smooth L1 loss $l(\hat{q_{i}},q_{i})$ with $\beta=0.01$ is calculated over the batch of target quaternions $Q\coloneqq\{q_{0},...,q_{n}\}$ and the predicted orientations $\hat{Q}\coloneqq\{\hat{q}_{0},...,\hat{q}_{n}\}$. The same procedure is repeated for the negated batch of target orientations $-Q$, so that $l(\hat{q_{i}},-q_{i})$ is calculated. The resulting error batches of shape $(N\times 4)$ are summed up row-wise. Then the minimum between the distance with the positive $q_{i}$ and the negative target rotation $-q_{i}$ is chosen for every index $i$. Finally, the batch mean is calculated over the resulting batch of minimum sums of errors.
@@ III-F Motion Planning
Grasping motions can quickly be generated by a small set of IK queries that are dynamically adapted towards a target position, while always maintaining the same orientation[34, 6]. However, the approach lacks scalability to other robot workspaces and can lead to unnatural movements that, in the worst case, include dynamic singularities when the velocity and acceleration profiles of the physical motors are insufficiently tuned. Our approach utilizes Bézier curves for a motion planner. They are traditionally a popular motion planning method for Cartesian robots, e.g. sorting robots[35]. The Bézier procedure, according to[36], interpolates a smooth trajectory in Cartesian space between a set of $n$ control points $C$, where $c_{0}$ is the start point and $c_{n}$ is the target point, as defined in Eq. 10:
@@ IV-A Hyperparameter Optimization
The hyperparameters of our MLP models are optimized with the Optuna [39] framework. $200$ different hyperparameter configurations are evaluated for each robot. The configurations are sampled with a Tree-structured Partizan Estimator (TPE) and a hyper-band pruner that stops unsuccessful trials which are likely to show low performance early. We optimize the following hyperparameters: the batch size, the learning rate (Lr), the number of layers, the number of neurons per hidden layer, the number of tanh (No. tanh) layers at the end of the network, the position weight ($w_{pos}$), and the orientation weight ($w_{rot}$). The parameter configurations of the most successful runs for each robot are shown in Table I.
@@
Robot pos.a $\sigma_{\hat{p}}$ rot.b $\sigma_{\hat{q}}$ success timec pos.a $\sigma_{\hat{p}}$ rot.b $\sigma_{\hat{q}}$ success timec pos.a $\sigma_{\hat{p}}$ rot.b $\sigma_{\hat{q}}$ success timec
@@
TABLE III: The precision of CycleIK compared to SOTA neuro-generative IK solvers IKFlow and EEM. The positional error (pos.) and rotational error (rot.) are reported in $\mathrm{mm}$ and and the runtime (time) on a batch of $100$ poses in $\mathrm{ms}$.
@@
Robot pos. rot. time pos rot time pos rot.
@@ IV-B Precision
The neuro-inspired inverse kinematics method CycleIK is evaluated on five different robot platforms in simulation: the NICOL[6], NICO[8], NASA Valkyrie[16], Franka Panda[15], and Fetch[17] robot. We evaluate the genetic algorithm BioIK[13] and the Jacobian-based Trac-IK[14] on the same test data sets for a baseline. Both numerical IK algorithms are available as a plug-in in Moveit[12]. The results of our experiments can be seen in Table II. Our test data sets contain $200,000$ poses for each robot that were randomly sampled in the joint space of the robots with Moveit (See Sec. III-D).
@@
Furthermore, CycleIK is evaluated against the state-of-the-art neuro-generative IK methods IKFlow[19] and the Euclidean Equivariant Models (EEMs) proposed by Limoyo et al.[18]. The results are shown in Table III. The experimental results for IKFlow and EEM are taken from the corresponding publication. The mean positional and mean rotational error are the only metrics reported in both other papers. Limoyo et al. do not report the inference time, which is likely high as their approach creates multiple embeddings via GNNs in every query. IKFlow reports the average runtime over batches that contain $100$ poses each. We adopted our experiment to this setup and always inferred the IK solution for a batch that holds $100$ poses, which results in a slightly higher runtime compared to the results that are reported in Table II, but the same positional and rotational precision as the MLP models deterministically return the same solution for the identical pose, independent of the batch size during inference.
CycleIK shows the overall highest precision and is only outperformed by the EEM approach regarding the rotational error on the Franka Panda robot test set. It has to be noted that IKFlow and EEMs are generative approaches that are capable of sampling the pose null space, while the CycleIK MLP is trained to return exactly one solution for every pose, conditioned to the secondary kinematic constraints that were set during the offline training. Consequently, it has to be acknowledged that the complexity of the task is thus slightly relieved for the CycleIK MLP when compared to the IKFlow and EEM approaches. The increase of the batch size from $1$ to $100$ during inference only had a small effect on the average inference time per batch, the runtime for the Franka Panda robot increased by $0.02\,\mathrm{ms}$ to $0.41\,\mathrm{ms}$, and for the Valkyrie robot by $0.05\,\mathrm{ms}$ to $0.43\,\mathrm{ms}$. Overall, batch sizes up to $1000$ on average did not increase the runtime above $1\,\mathrm{ms}$ in our preliminary tests.
@@ IV-C Grasping
The results of our experiment can be seen in Fig. 5. We defined five different result classes into which every grasp approach is classified during the experiment: Success, Slipped, Not Grasped, Detection, and LLM. A grasp approach is considered Success when the object is lifted from the table and correctly disposed of into the bin at the outer workspace. An approach is considered Slipped when it is correctly lifted from the table vertically but slips out of the hand during the transfer to the bin. Approaches that fail to grasp the object correctly and that are not able to lift it from the table are classified as Not Grasped. The Detection class contains approaches that cannot successfully be executed because the object detector predicts a wrong or no class at all, or a completely wrong location. The LLM class contains approaches in which the language model is aware of all the objects on the table but does not execute the grasp action on all of them.
@@ V Conclusions
This work successfully introduced a novel Bézier-based non-collision-free motion planner that is based on the neuro-inspired inverse kinematics solver CycleIK. The motion planner was deployed to the physical humanoid robots NICO and NICOL in an embodied agent setup. Our method reached a success rate of $72\%$ on the NICOL robot and $82\%$ on the NICO robot in a human-in-the-loop grasping scenario. New loss functions for the positional and rotational error calculation were introduced to CycleIK that were derived from the smooth L1 loss. We have shown that the new CycleIK method can compete with and partially outperforms well-known numerical approaches such as BioIK and Trac-IK, but also state-of-the-art neuro-inspired approaches like IKFlow in terms of precision. The best iterative numerical approach performs slightly better than CycleIK but it requires a longer runtime. However, CycleIK provides an ideal choice, when runtime is a major critical factor and performs very good for the NICO and NICOL robots. We have shown that CycleIK is one of the fastest batched IK methods available and offers quick planning of robot motions in Cartesian space.
@@ Acknowledgement
The authors thank Erik Strahl and Matthias Kerzel for the extensive discussions and thought-provoking impulses on inverse kinematics during the development of CycleIK. The authors thank Ozan Özdemir for his efforts in the development of the camera model of the NICO robot. The authors thank Hassan Ali for his contributions to the implementation of ViLD on the NICOL robot.

### v2

@@ Abstract
This paper introduces a novel zero-shot motion planning method that allows users to quickly design smooth robot motions in Cartesian space. A Bézier curve-based Cartesian plan is transformed into a joint space trajectory by our neuro-inspired inverse kinematics (IK) method CycleIK, for which we enable platform independence by scaling it to arbitrary robot designs. The motion planner is evaluated on the physical hardware of the two humanoid robots NICO and NICOL in a human-in-the-loop grasping scenario. Our method is deployed with an embodied agent that is a large language model (LLM) at its core. We generalize the embodied agent, that was introduced for NICOL, to also embody NICO. The agent can execute a discrete set of physical actions and allows the user to verbally instruct various different robots. We contribute a grasping primitive to its action space that allows for precise manipulation of household objects. The updated CycleIK [^2] Available on GitHub: https://github.com/jangerritha/cycleik method is compared to popular numerical IK solvers and state-of-the-art neural IK methods in simulation and is shown to be competitive with or outperform all evaluated methods when the algorithm runtime is very short. The grasping primitive is evaluated on both NICOL and NICO robots with a reported grasp success of 72% to 82% for each robot, respectively.
© 2024 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
@@ I Introduction
Precise inverse kinematics is the key factor for successful grasping motions in many robotic manipulation applications. IK methods predict joint angles that align the robot end effector with a given position and orientation. They therefore collapse the action space of higher layered motion planners to the 6D pose of the end effector, while the alternative is to control the robot directly in its joint space which poorly scales for robots with varying degrees of freedom (DoF). The complexity of the IK task increases drastically for redundant kinematic chains with more than six DoF, since for non-redundant chains a specific pose maps exactly to one joint configuration, while for redundant manipulators infinite solutions generally exist, which makes the problem ambiguous. The number of neuro-inspired IK methods that are precise enough to be applied to physical hardware is increasing but there are only a few approaches that utilize them for neural motion planning[1, 2]. Many motion planners exist that use neural networks to sample the robot joint space and deploy classical planning algorithms like MPNet[3]. Thus, only the runtime of the sampling decreases, while the latency for the motion planning itself stays high. Other approaches plan directly in Cartesian space and transform the plan to the joint space via inverse kinematics such as CppFlow[1]. Neural networks that learn to solve the IK task, which allows for parallel batched queries, can dramatically accelerate Cartesian planning methods but decrease the precision. Numerical IK approaches in difference have to solve each pose in an individual optimization run, resulting in a trade-off between precision and runtime, for which we propose one of the most runtime-efficient motion planning systems available.
@@
The neural inverse kinematics solver CycleIK[4] is the central component of our proposed motion planning approach. CycleIK is scaled to arbitrary robot designs by a replacement of the loss functions for the positional and rotational error with two new metrics that are derived from the Smooth L1 loss[5]. The new loss functions enable high precision within ten epochs of training, while our former publication suggested an ideal training time of $100$ to $300$ epochs using the old loss functions. The low-level motion planner interpolates a Bézier curve between the current pose and a target pose and then converts the Cartesian plan into a joint space trajectory in a single step via CycleIK. Our method is deployed with the high-level embodied agent that is introduced to the NICOL[6] robot by Allgeuer et al.[7]. We generalize the embodied agent architecture over our humanoid robot platform to also embody the NICO[8] robot.
@@
The updated CycleIK methodology is evaluated against two well-established IK methods available in Moveit[12], the genetic algorithm BioIK[13] and the Jacobian-based method Trac-IK[14]. The algorithms are compared on five different robot platforms: NICO, NICOL, Franka Emika Panda[15], NASA Valkyrie[16], and the Fetch[17] robot. In addition, the proposed IK method is compared against two state-of-the-art neuro-generative IK approaches by Limoyo et al.[18] and Ames et al.[19]. Finally, the grasp success of our method is evaluated on the NICO and the NICOL robots, for $100$ grasp attempts on each robot.
@@ III Method
The presented method fundamentally updates our neuro-inspired IK approach CycleIK[4] to enable platform independence for our method. Most importantly, the loss functions for the positional and rotational errors are exchanged by new Smooth L1-based metrics. The CycleIK multi-layer perceptron (MLP) is a single-solution IK solver that deterministically predicts a single vector of joint angles for a specific pose. Secondary constraints, such as zero-joint goals or pose goals for other links in the kinematic chain, can optionally be applied during the training to influence the kinematic behavior of the model. CycleIK also provides a generative adversarial network that is capable of sampling the null space of a specific pose for redundant kinematic chains with higher than six DoF. The GAN can enable obstacle-avoiding motion planning, but as this study does not consider obstacle-avoidance, the MLP is sufficient since it is more applicable and will therefore be used to solve the inverse kinematics problem in our experiments.
@@
Figure 3: Overview of the CycleIK training algorithm. Optional secondary kinematic constraints can either be set in the joint space or Cartesian space.
@@ III-B Training
We utilize the geometric forward kinematics solver PyTorch Kinematics[30] that allows for batched FK queries as it calculates the end effector pose for a particular joint state by matrix multiplications and allows for computational acceleration by utilizing the GPU via PyTorch. The operations used by PyTorch Kinematics are differentiable and can therefore be used during the network training without blocking the propagation of the gradient. Every training iteration processes a batch of poses with a minimum size of $100$, while the exact batch size is a hyperparameter that is optimized per robot. The learning rate is linearly decreased to zero over the training time.
@@ III-C Metrics
$l(a,b)=\begin{cases}\tfrac{1}{2}(a-b)^{2}/\beta,&\text{if $|a-b|<\beta$}\\ |a-b|-\tfrac{1}{2}\beta,&\text{otherwise}\end{cases}$ (5)
@@
First, the element-wise smooth L1 loss $l(\hat{q_{i}},q_{i})$, with $\beta=0.01$ corresponding to $\sim 1^{\circ}$, is calculated over the batch of target quaternions $Q\coloneqq\{q_{0},...,q_{n}\}$ and the predicted orientations $\hat{Q}\coloneqq\{\hat{q}_{0},...,\hat{q}_{n}\}$. The same procedure is repeated for the negated target orientations $-Q$ so that $l(\hat{q_{i}},-q_{i})$ is calculated. The resulting errors of shape $(N\times 4)$ are summed up row-wise. Then the minimum between the distance to the positive $q_{i}$ and the negative target rotation $-q_{i}$ is chosen for every index $i$. Finally, the batch mean is calculated over the resulting batch of minimum sums of errors.
@@ III-F Motion Planning
Grasping motions can quickly be generated by a small set of IK queries that are dynamically adapted towards a target position, while always maintaining the same orientation[34, 6], [35]. However, the approach lacks scalability to other robot workspaces and can lead to unnatural movements that, in the worst case, include dynamic singularities when the velocity and acceleration profiles of the physical motors are insufficiently tuned. Our approach utilizes Bézier curves for a motion planner. They are traditionally a popular motion planning method for Cartesian robots, e.g. sorting robots[36]. The Bézier procedure, according to[37], interpolates a smooth trajectory in Cartesian space between a set of $n$ control points $C$, where $c_{0}$ is the start point and $c_{n}$ is the target point, as defined in Eq. 10:
@@ IV-A Hyperparameter Optimization
The hyperparameters of our MLP models are optimized with the Optuna [40] framework. $200$ different hyperparameter configurations, that are sampled with a Tree-structured Partizan Estimator (TPE), are evaluated for each robot. We optimize the following hyperparameters: batch size, learning rate (Lr), number of layers, number of neurons per hidden layer, number of tanh (No. tanh) layers at the end of the network, position weight ($w_{pos}$), and orientation weight ($w_{rot}$). The evaluation of each configuration is conducted with a fixed weighting of the losses with a factor of $10$ for $w_{pos}$ and $1$ for $w_{rot}$. The parameter configurations of the most successful runs for each robot are shown in Table I.
@@
Robot pos.a $\sigma_{pos}$ rot.b $\sigma_{rot}$ success timec pos.a $\sigma_{pos}$ rot.b $\sigma_{rot}$ success timec pos.a $\sigma_{pos}$ rot.b $\sigma_{rot}$ success timec
@@
TABLE III: Overview of the precision of SOTA neuro-inspired IK solvers CycleIK, IKFlow, and EEM. The positional error (pos.) and rotational error (rot.) are reported in $\mathrm{mm}$ and and the runtime (time) on a batch of $100$ poses in $\mathrm{ms}$.
@@
Robot pos. rot. time pos. rot. time pos. rot.
@@ IV-B Precision
The neuro-inspired inverse kinematics method CycleIK is evaluated on five different robot platforms in simulation: the NICOL[6], NICO[8], NASA Valkyrie[16], Franka Panda[15], and Fetch[17] robot. No secondary constraints were set for the training. We evaluate the genetic algorithm BioIK[13] and the Jacobian-based Trac-IK[14] on the same test data sets for a baseline. Both numerical IK algorithms are available as a plug-in in Moveit[12]. The results of our experiments can be seen in Table II. Our test data sets contain $200,000$ poses for each robot that were randomly sampled in the joint space of the robots with Moveit (See Sec. III-D).
@@
Furthermore, Table III provides an overview of state-of-the-art neuro-inspired IK methods, including IKFlow[19] and the Euclidean Equivariant Models (EEMs) proposed by Limoyo et al.[18]. The experimental results for IKFlow and EEM are taken from the corresponding publication. The mean positional and mean rotational error are the only metrics reported in both other papers. Limoyo et al. do not report the inference time. IKFlow reports the average runtime over batches that contain $100$ poses each. We adopted our experiment to this setup and always inferred the IK solution for a batch that holds $100$ poses, which results in a slightly higher runtime compared to the results that are reported in Table II, but the same positional and rotational precision as the MLP models deterministically return the same solution for the identical pose, independent of the batch size.
CycleIK shows the overall highest precision and is only outperformed by the EEM approach regarding the rotational error on the Franka Panda robot test set. It has to be noted that IKFlow and EEMs are generative approaches that are capable of sampling the pose null space, while the CycleIK MLP deterministically returns one solution for every pose. In addition, EEMs are a multi-robot IK method that is capable of solving the IK problem for five different robots with a single model. Consequently, it has to be acknowledged that the complexity of the task is thus strongly relieved for the CycleIK MLP when compared to the IKFlow and EEM approaches. The increase of the batch size from $1$ to $100$ during inference only had a small effect on the average inference time per batch, the runtime for the Franka Panda robot increased by $0.02\,\mathrm{ms}$ to $0.41\,\mathrm{ms}$, and for the Valkyrie robot by $0.05\,\mathrm{ms}$ to $0.43\,\mathrm{ms}$. Overall, batch sizes up to $1000$ on average did not increase the runtime above $1\,\mathrm{ms}$ in our preliminary tests. While a direct comparison is complex, all three models present a new class of neural IK methods that reach a millimeter range of error compared to centimeters as in many previous approaches in this field, e.g. [21] or [22].
@@ IV-C Grasping
The results of our experiment can be seen in Fig. 5. We defined five different result classes into which every grasp approach is classified during the experiment: Success, Slipped, Not Grasped, Detection, and LLM. A grasp approach is considered Success when the object is lifted from the table and correctly disposed of into one of the bins. An approach is considered Slipped when it is correctly lifted from the table vertically but slips out of the hand during the transfer to the bin. Approaches that fail to grasp the object correctly and that are not able to lift it from the table are classified as Not Grasped. The Detection class contains approaches that cannot successfully be executed because the object detector predicts a wrong or no class at all, or a completely wrong location. The LLM class contains approaches in which the language model is aware of all the objects on the table but does not execute the grasp action on all of them.
@@ V Conclusions
This work successfully introduced a novel Bézier-based non-collision-free motion planner that is based on the neuro-inspired inverse kinematics solver CycleIK. New loss functions for the positional and rotational error calculation were introduced to CycleIK that were derived from the smooth L1 loss and enable high precision within a short training time. We have shown that the new CycleIK method can compete with and partially outperforms well-known numerical approaches such as BioIK and Trac-IK, but also state-of-the-art neuro-inspired approaches like IKFlow in terms of precision. We have shown that CycleIK is one of the fastest batched IK methods available and offers quick planning of robot motions in Cartesian space. The best iterative numerical approach performs better than CycleIK but it requires a longer runtime. However, CycleIK provides an ideal choice, when the runtime is a major critical factor and performs very good for the NICO and NICOL robots. The motion planner was deployed to the physical humanoid robots NICO and NICOL in an embodied agent setup. Our method reached a success rate of $72\%$ on the NICOL robot and $82\%$ on the NICO robot in a human-in-the-loop grasping scenario.
@@ Acknowledgement
The authors thank Erik Strahl and Matthias Kerzel for discussions on inverse kinematics for CycleIK; Ozan Özdemir for camera development of the NICO robot; Hassan Ali for his contributions to the object detection on the NICOL robot.